# Production and decay of hyperons in transversely polarized electron-positron collider

### v1

@@ I Introduction
Transversely-polarized beams at an electron-positron collider are particularly interesting in the search for new sources of CP violation through the measurement of CP odd azimuthal asymmetry Ananthanarayan and Rindani (2004). They also offer a potential opportunity for revealing fundamental interactions and probing new physics Moortgat-Pick et al. (2008). For instance, the azimuthal modulation distributions would originate from the interference of operators of new physics beyond the Standard Model (SM) with those in the SM Wen et al. (2023). Additionally, transverse-polarization monitoring is a key technique for beam energy calibration using the resonant depolarization method, thereby improving the measurement of the mass and width of narrow resonances, such as the $Z-$boson and $J/\psi$ Chen et al. (2022).
The well-known Sokolov-Ternov effect Sokolov and Ternov (1963) refers to the spin-flip processes through synchrotron radiation emission, which leads to an increase in polarization effect with the rise in energy. This results in a natural build-up of transverse polarization through the competition between radiation self-polarization and spin diffusion Baier and Fadin (1965); Jackson (1976). In the mid-1970s, the ACO storage ring at Orsay Montague (1984) and the improved VEPP-2M at the Budker Institute of Nuclear Physics (BINP) Kurdadze et al. (1976) both approached the theoretical limit of polarization levels, approximately ${8}/{5\sqrt{3}}\simeq$ 92.4%. At a beam energy of 3.7 GeV, SPEAR II found the equilibrium value of polarization to be about 0.76 Learned et al. (1975). The degree of transverse beam polarization at the CERN Large Electron Positron storage ring (LEP) was observed to be around 9.1% Knudsen et al. (1991). In the future, it is possible to polarize the beams up to 10% in a few hours at a designed energy of 45.6 GeV at the high-energy Circular Electron-Positron Collider (CEPC) Nikitin (2019). The same level of transverse polarization of the beams is expected in the Future Circular Collider (FCC-ee) Blondel et al. (2019). As the first and only ring to provide longitudinal lepton spin polarization at high energy (27.5 GeV), the production of polarized electron and positron beams in HERA relied on the Sokolov-Ternov effect, where the spin rotators on either side of the interaction points converted the polarization of the beam from transverse to longitudinal. High luminosity at SuperKEKB prevents the use of the Sokolov-Ternov effect from yielding longitudinal polarization Asner et al. (2022).
@@
where the Lorentz factor is $\gamma=E_{e}/m_{e}$, leading to $\tau_{0}=2.8$ hours at $E_{e}=2.0$ GeV, using an average radius $R=38$ meters and an effective radius $\rho=9.3$ meters for BEPCII (Beijing Electron-Positron Collider). The depolarization resonances from imperfections in the magnetic field occur at certain energies determined by gyromagnetic anomaly, leading into vanishing polarization of beams. The realistic situation is, of course, more complicated due to many other sources of depolarization resonances. However, the degree of transverse polarization of the beams at BEPCII and future super $\tau$-charm facility (STCF) Achasov et al. (2024), away from the depolarization resonances, is expected to be of fair magnitude, resulting in a sizeable impact on the angular distribution of final particles.
For a long time, spin asymmetries and correlations in hyperon-pair production in unpolarized electron-positron collisions have been proposed to measure simultaneously the electric-magnetic form factors and decay parameters of hyperons and anti-hyperons Czyz et al. (2007); Chen and Ping (2007). This approach allows for the investigation of CP nonconservation parameters with high precision Lee and Yang (1957); Donoghue and Pakvasa (1985); Donoghue et al. (1986). Recently, CP violation in hyperon decays at super-charm-tau factories with a longitudinally polarized electron beam has been investigated Salone et al. (2022); Zeng et al. (2023), building upon earlier efforts Dubnickova et al. (1996); Brodsky et al. (2004); Tomasi-Gustafsson et al. (2005). However, the effects from transverse beam polarization were never taken into account in the analysis and simulation of the $e^{+}e^{-}\to B\overline{B}$ reactions, except for some earlier attempts on angular distributions Tsai (1975); Bletzacker and Nieh (1976); Hikasa (1986). In particular, the hyperon transverse polarization in $\psi(3686)\rightarrow\Lambda\bar{\Lambda}$ Ablikim et al. (2023a) and $\Sigma^{+}\bar{\Sigma}^{-}$ decays Ablikim et al. (2020) are expected to be affected by the transverse polarizations of the lepton beams, as is the case for $\psi(3770)$ Ablikim et al. (2022a). The transverse polarization of double-strange baryons $\Xi$ observed with unprecedented accuracy by the BESIII collaboration in $\psi(3686)\rightarrow\Xi^{-}\bar{\Xi}^{+}$ Ablikim et al. (2022b), $\Xi^{0}\bar{\Xi}^{0}$ Ablikim et al. (2023b), $\Xi(1530)^{-}\bar{\Xi}(1530)^{+}$ and $\Xi(1530)^{-}\bar{\Xi}^{+}$ Ablikim et al. (2019a) also requires further scrutiny. The polarization of most strange baryons $\Omega$ in $\psi(3686)\to\Omega^{-}\bar{\Omega}^{+}$ Ablikim et al. (2021) deserves more attention if enough events are accumulated. Additionally, data from different isospin channels would be helpful for understanding the hyperon electromagnetic form factors Dai et al. (2023).
@@ II.0.1 Spin Density Matrix
Due to synchrotron radiation when positrons and electrons circulate in the storage ring, the transition probabilities of the two spin projections of positrons and electrons, guided by the magnetic field in the storage ring, are different. This causes the spin orientation of positrons to tend towards the direction of the guiding magnetic field, while the orientation of electrons is opposite. Consequently, as the lepton beam remains in the storage ring for an extended period, they will acquire a transverse polarization $\mathcal{P}_{t}=p_{x}+ip_{y}=P_{T}e^{i\phi_{+}}$ and $\bar{\mathcal{P}}_{t}=\bar{p}_{x}+i\bar{p}_{y}=\bar{P}_{T}e^{i\phi_{-}}$ for positron and electron, respectively, where $P_{T}=|\mathcal{P}_{t}|$. Here, $p_{x}(\bar{p}_{x})$ represents the degree of transverse polarization in the scattering plane, and $p_{y}(\bar{p}_{y})$ represents the degree of polarization perpendicular to the scattering plane Moortgat-Pick et al. (2008). The angles $\phi_{+}$ and $\phi_{-}$ represent the azimuthal angles of the positron and electron polarizations, respectively, with respect to the lab system. In a positron-electron annihilation experiment with symmetric beam energy, the Sokolov–Ternov effect requires the equal degree of polarization $P_{T}=\bar{P}_{T}$ and $\phi_{+}=\phi_{-}=\pi/2$. This means that positron and electron have the same polarization vector in the individual helicity frame  [^1] In an $e^{+}e^{-}$ storage ring, the $z$-axes of the helicity frames for the electron and positron are aligned with their respective directions of motion, and they share the same $x$-axis along the radical direction, while the vertical $y$-axis is oriented in the opposite direction.. The spin density matrix of the leptons is represented in their helicity frame as:
@@
$\rho^{\gamma^{*}/\psi}$ $=$ $\frac{1}{2}\left(\begin{array}[]{ccc}(1+\mathcal{P}_{z})(1-\bar{\mathcal{P}}_{z})&0&P_{T}^{2}\\[5.0pt] 0&0&0\\ P_{T}^{2}&0&(1-\mathcal{P}_{z})(1+\bar{\mathcal{P}}_{z})\\ \end{array}\right),$ (12)
@@
$\rho^{\gamma^{*}/\psi}$ $=$ $\frac{1}{3}\left(\begin{array}[]{ccc}1+{3q_{z}\over 2}+\sqrt{3\over 2}(T_{xx}+T_{yy}+2T_{zz})&{3(q_{x}-iq_{y})\over 2\sqrt{2}}&\sqrt{3\over 2}(T_{xx}-T_{yy})\\[5.0pt] {3(q_{x}+iq_{y})\over 2\sqrt{2}}&1+\sqrt{6}T_{xx}+\sqrt{6}T_{yy}&{3(q_{x}-iq_{y}))\over 2\sqrt{2}}\\ \sqrt{3\over 2}(T_{xx}-T_{yy})&{3(q_{x}+iq_{y}))\over 2\sqrt{2}}&1-{3q_{z}\over 2}+\sqrt{3\over 2}(T_{xx}+T_{yy}+2T_{zz})\\ \end{array}\right).$ (16)
Comparing the elements of the $\rho^{\psi}$ matrix, we can obtain $q_{x}=q_{y}=0,~{}q_{z}={2\over 3}\mathcal{P}_{z},~{}T_{xx}={3P_{t}^{2}-1\over 2\sqrt{6}},~{}T_{yy}=-{1+3P_{t}^{2}\over 2\sqrt{6}},~{}T_{zz}=1/\sqrt{6}$ with other $T_{ij}(i\neq j)=0$. It can be seen that the tensor polarization of $\rho^{\gamma^{*}/\psi}$ comes from the spin correlation and transverse polarization of the initial lepton beams, with linear polarization being solely from the longitudinal polarization of beams. In the following paper, we focus solely on the transverse polarization of the initial lepton beams. The study of a longitudinally polarized electron beam has been recently addressed independently in other studies Salone et al. (2022); Zeng et al. (2023).
@@ II.0.2 $e^{+}e^{-}\to\gamma^{*}/\psi\to\Lambda(p\pi^{-})\bar{\Lambda}(\overline{p}\pi^{+})$
In the laboratory system, the decayed $\Lambda$ particle moves in the direction defined by polar and azimuthal angles $(\theta,\phi)$. We calculate the joint angular distribution in the $\Lambda$ helicity system, as illustrated in Fig. 1. In this system, the $z$-axis aligns with the $\Lambda$ particle’s direction of motion, the $y$-axis is perpendicular to the $\Lambda$ production plane, i.e., $\hat{y}=\hat{p}_{+}\times\hat{p}_{\Lambda}$, where $\hat{p}_{+}$ and $\hat{p}_{\Lambda}$ are the unit momentum vectors for the positron and $\Lambda$, respectively. The $x$-axis lies in the $\Lambda$ production plane, forming a right-handed $x$-$y$-$z$ coordinate system. After boosting the momenta of $p$ and $\pi^{-}$ to the $\Lambda\to p\pi^{-}$ decay plane, they remain in the same plane. The angle between the proton and $\Lambda$ momentum is defined as $\theta_{1}$, and the angle between the $\Lambda$ production and decay plane is defined as $\phi_{1}$. Similar helicity angles $(\theta_{2},\phi_{2})$ are defined in the same manner.
@@
$\rho_{1}^{i,j}(\theta,\phi)$$\equiv$ $\sum\limits_{k=\pm 1}\rho^{\gamma^{*}/\psi}_{k,k^{\prime}}{\cal D}_{k,i}^{1*}(\phi,\theta,0)\,{\cal D}^{1}_{k^{\prime},j}(\phi,\theta,0)$ (17)
@@
$+$ $\frac{3P_{T}^{2}}{3+\alpha_{\psi}}\cdot\left(\begin{array}[]{cccc}\alpha_{\psi}\sin^{2}\!\theta\cos\!2\phi&-{\beta_{\psi}{\sin\!\theta}}\sin\!2\phi&-{\beta_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi&0\\ -{\beta_{\psi}{\sin\!\theta}}\sin\!2\phi&(\alpha_{\psi}+\cos^{2}\!\theta)\cos\!2\phi&-(1+\alpha_{\psi})\cos\!\theta\sin\!2\phi&-{\gamma_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi\\ {\beta_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi&(1+\alpha_{\psi})\cos\!\theta\sin\!2\phi&(1+\alpha_{\psi}\cos\!\theta)\cos\!2\phi&-{\gamma_{\psi}{\sin\!\theta}}\sin\!2\phi\\ 0&{\gamma_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi&-{\gamma_{\psi}{\sin\!\theta\sin\!2\phi}}&-\sin^{2}\!\theta\cos\!2\phi\\ \end{array}\right),$ (37)
@@
without longitudinal polarization component. Besides, transverse polarization of beams introduce two new correlations $C_{xy}=-C_{yx}$, $C_{yz}=C_{zy}$ in addition to the four existing ones, which further constrain the decay parameters of hyperons, and thus provides more stringent tests of the CP violation.
@@
and Fig. 2 shows its three dimensional distribution. So that if the transverse polarization of the final-state particles is not measured, the effects of transverse polarizations are absent in the $\phi$-averaged cross section, albeit, presnt in the $\theta$-averaged cross section:
@@
with $(\theta_{0},\pi-\theta_{0})$ being the detector coverage of the solid angle around the interaction point (IP). Therefore the degree of transverse polarization would be measured through $e^{+}e^{-}\to\mu^{+}\mu^{-}$ and $e^{+}e^{-}$ azimuth angular distributions. The distribution of scattering angle has been previously explored in BES Bai et al. (1995), BESIII Ablikim et al. (2019b) and KEDR at the VEPP-4M Anashin et al. (2018), but the azimuth distributions were not received any attention.
There are five global parameters to describe a process $e^{+}e^{-}\to B\overline{B}$ followed by single-step weak two-body decays of the hyperon $B$ and the antihyperon $\overline{B}$ Perotti et al. (2019). For decay $B\!\to\!b\pi$ and the corresponding charge congugate (c.c.) decay mode $\overline{B}\!\to\!\overline{b}\overline{\pi}$, like $e^{+}e^{-}\to\gamma^{*}/\psi\to\Lambda(p\pi^{-})\overline{\Lambda}(\overline{p}\pi^{+})$, they are represented by the vector $\boldsymbol{\omega}\equiv(\alpha_{\psi},\Delta\Phi,\alpha_{-},{\alpha_{+}})$ with a priori known $P_{T}^{2}$ and $\alpha_{-}$ (or ${\alpha_{+}}$) being the decay parameter of $B\!\to\!b\pi$ (or $\overline{B}\!\to\!\overline{b}\overline{\pi}$). The joint angular distribution $\mathcal{W}(\boldsymbol{\xi})$ can be expressed with respect to the vector $\boldsymbol{\xi}\equiv(\Omega_{B},\Omega_{b},\Omega_{\overline{b}})$ representing a complete set of the kinematic variables describing a single-event configuration in the six dimensional phase space Hong et al. (2023); Zeng et al. (2023):
@@
Refer to captionRefer to caption
Figure 3: The transverse polarization $P_{y}$ and $P_{x}$ of $\Lambda$ hyperon in all azimuthal angle in comparism of upper and lower sphere. The parameters of $e^{+}e^{-}\to\psi(3686)\to\Lambda\overline{\Lambda}$ with $\alpha=0.69$ and $\Delta\Phi=23^{\circ}$ Ablikim et al. (2023a) are used.
@@
the latter of which is corresponding to the number of events in the range of upper semi-sphere $45^{\circ}$ to $135^{\circ}$ and $225^{\circ}$ to $315^{\circ}$, and the former is the number of events in the remaining lower semi-sphere. Then $\mathcal{W}^{+}(\xi)-\mathcal{W}^{-}(\xi)$ is propotional to those terms of $P_{T}^{2}\cos 2\phi$ dependence and $\mathcal{W}^{+}(\xi)+\mathcal{W}^{-}(\xi)$ integrates out the $P_{T}^{2}$ terms which are of $\cos 2\phi$ modulation. To select the $\sin 2\phi$ modulation:
@@
the former of which is corresponding to the number of events in the range of upper semi-sphere $0^{\circ}$ to $90^{\circ}$ and $180^{\circ}$ to $270^{\circ}$, and the latter is the number of events in the remaining lower semi-sphere. Fig. 3 shows the transverse polarization $P_{y}$ and $P_{x}$ of baryon in all azimuthal angle in comparism of upper and lower sphere. Instead, one can investigate the moments of the joint angular distribution as shown in Sec. III.

### v2

# Production and decay of hyperons in a transversely polarized electron-positron collider

@@ I Introduction
Transversely-polarized beams at an electron-positron collider are particularly interesting in the search for new sources of CP violation through the measurement of CP odd azimuthal asymmetry Ananthanarayan and Rindani (2004). They also offer a potential opportunity for revealing fundamental interactions and probing new physics Moortgat-Pick et al. (2008). For instance, some of specific modulations of azimuthal distributions would originate from the interference of operators of new physics beyond the Standard Model (SM) with those in the SM Wen et al. (2023). Additionally, transverse-polarization monitoring is a key technique for beam energy calibration using the resonant depolarization method, thereby improving the measurement of the mass and width of narrow resonances, such as the $Z-$ boson and $J/\psi$ Chen et al. (2022).
The well-known Sokolov-Ternov effect Sokolov and Ternov (1963) refers to the spin-flip processes through synchrotron radiation emission, which results in a natural build-up of transverse polarization through the competition between radiation self-polarization and spin diffusion Baier and Fadin (1965); Jackson (1976). In the mid-1970s, the ACO storage ring at Orsay Montague (1984) and the improved VEPP-2M at the Budker Institute of Nuclear Physics (BINP) Kurdadze et al. (1976) both approached the theoretical limit of polarization levels, approximately $P_{0}={8}/{5\sqrt{3}}\simeq$ 92.4%. At a beam energy of 3.7 GeV, SPEAR II found the equilibrium value of polarization to be about 0.76 Learned et al. (1975). The degree of transverse beam polarization at the CERN Large Electron Positron storage ring (LEP) was observed to be around 9.1% Knudsen et al. (1991). In the future, it is possible to polarize the beams up to 10% in a few hours at a designed energy of 45.6 GeV at the high-energy Circular Electron-Positron Collider (CEPC) Nikitin (2019). The same level of transverse polarization of the beams is expected in the Future Circular Collider (FCC-ee) Blondel et al. (2019). As the first and only ring to provide longitudinal lepton spin polarization at high energy (27.5 GeV), the production of polarized electron and positron beams in HERA relied on the Sokolov-Ternov effect, where the spin rotators on either side of the interaction points converted the polarization of the beam from transverse to longitudinal, or vise versa. High luminosity at SuperKEKB prevents the use of the Sokolov-Ternov effect from accumulating or yielding longitudinal polarization Asner et al. (2022).
@@
where the Lorentz factor is $\gamma=E_{e}/m_{e}$, leading to $\tau_{0}=2.8$ hours at $E_{e}=2.0$ GeV, using an average radius $R=38$ meters and an effective radius $\rho=9.3$ meters for BEPCII (Beijing Electron-Positron Collider). The degree of transverse polarization $P_{T}=P_{0}(1-e^{-t/\tau_{0}})$ reaches 0.28 after one hour of beam injection, and it increases with the rise of beam energy over the same duration of injection time at the same accelerator. The depolarization resonances from imperfections in the magnetic field occur at certain energies determined by gyromagnetic anomaly, leading into vanishing polarization of beams. The realistic situation is, of course, more complicated due to many other sources of depolarization resonances. However, the degree of transverse polarization of the beams at BEPCII and future super $\tau$-charm facility (STCF) Achasov et al. (2024), away from the depolarization resonances, is expected to be of fair magnitude, resulting in a sizable impact on the angular distribution of final particles.
For a long time, spin asymmetries and correlations in hyperon-pair production in unpolarized electron-positron collisions have been proposed to measure simultaneously the electric-magnetic form factors and decay parameters of hyperons and anti-hyperons Czyz et al. (2007); Chen and Ping (2007); Fäldt and Kupsc (2017). This approach allows for the investigation of CP nonconservation parameters with high precision Lee and Yang (1957); Donoghue and Pakvasa (1985); Donoghue et al. (1986). Recently, CP violation in hyperon decays at super-charm-tau factories with a longitudinally polarized electron beam has been investigated Salone et al. (2022); Zeng et al. (2023), building upon earlier efforts Dubnickova et al. (1996); Brodsky et al. (2004); Tomasi-Gustafsson et al. (2005). However, the effects from transverse beam polarization were never taken into account in the analysis and simulation of the $e^{+}e^{-}\to B\overline{B}$ reactions, except for some earlier attempts on angular distributions Tsai (1975); Bletzacker and Nieh (1976); Hikasa (1986). In particular, the hyperon transverse polarization in $\psi(3686)\rightarrow\Lambda\bar{\Lambda}$ Ablikim et al. (2023a) and $\Sigma^{+}\bar{\Sigma}^{-}$ decays Ablikim et al. (2020) is expected to be affected by the transverse polarization of the lepton beams, as is the case for $\psi(3770)$ Ablikim et al. (2022a). The transverse polarization of double-strange baryons $\Xi$ observed with unprecedented accuracy by the BESIII collaboration in $\psi(3686)\rightarrow\Xi^{-}\bar{\Xi}^{+}$ Ablikim et al. (2022b); Liu et al. (2023), $\Xi^{0}\bar{\Xi}^{0}$ Ablikim et al. (2023b), $\Xi(1530)^{-}\bar{\Xi}(1530)^{+}$ and $\Xi(1530)^{-}\bar{\Xi}^{+}$ Ablikim et al. (2019a) also requires further scrutiny. The polarization of most strange baryons $\Omega$ in $\psi(3686)\to\Omega^{-}\bar{\Omega}^{+}$ Ablikim et al. (2021) deserves more attention if enough events are accumulated. Additionally, data from different isospin channels would be helpful for understanding the hyperon electromagnetic form factors Dai et al. (2023).
@@ II.0.1 Spin Density Matrix
Due to synchrotron radiation when positrons and electrons circulate in the storage ring, the transition probabilities of the two spin projections of positrons and electrons, guided by the magnetic field in the storage ring, are different. This causes the spin orientation of positrons to tend toward the direction of the guiding magnetic field, while the orientation of electrons is opposite. Consequently, as the lepton beam remains in the storage ring for an extended period, they will acquire a transverse polarization $\mathcal{P}_{t}=p_{x}+ip_{y}=P_{T}e^{i\phi_{+}}$ and $\bar{\mathcal{P}}_{t}=\bar{p}_{x}+i\bar{p}_{y}=\bar{P}_{T}e^{i\phi_{-}}$ for positron and electron, respectively, where $P_{T}=|\mathcal{P}_{t}|$. Here, $p_{x}(\bar{p}_{x})$ represents the degree of transverse polarization in the scattering plane, and $p_{y}(\bar{p}_{y})$ represents the degree of polarization perpendicular to the scattering plane Moortgat-Pick et al. (2008). The angles $\phi_{+}$ and $\phi_{-}$ represent the azimuthal angles of the positron and electron polarizations, respectively, with respect to the lab system. In a positron-electron annihilation experiment with symmetric beam energy, the Sokolov–Ternov effect requires the equal degree of polarization $P_{T}=\bar{P}_{T}$ and $\phi_{+}=\phi_{-}=\pi/2$. This means that the positron and the electron have the same polarization vector in the individual helicity frame  [^1] In an $e^{+}e^{-}$ storage ring, the $z$-axes of the helicity frames for the electron and positron are aligned with their respective directions of motion, and they share the same $x$-axis along the radical direction, while the vertical $y$-axis is oriented in the opposite direction.. The spin density matrix of the leptons is represented in their helicity frame as:
@@
$\rho^{\gamma^{*}/\psi}$ $=$ $\frac{1}{2}\left(\begin{array}[]{ccc}(1-\mathcal{P}_{z})(1+\bar{\mathcal{P}}_{z})&0&P_{T}^{2}\\[5.0pt] 0&0&0\\ P_{T}^{2}&0&(1-\mathcal{P}_{z})(1+\bar{\mathcal{P}}_{z})\\ \end{array}\right),$ (12)
@@
$\rho^{\gamma^{*}/\psi}$ $=$ $\frac{1}{3}\left(\begin{array}[]{ccc}1+{3q_{z}\over 2}+\sqrt{3\over 2}(T_{xx}+T_{yy}+2T_{zz})&{3(q_{x}-iq_{y})\over 2\sqrt{2}}&\sqrt{3\over 2}(T_{xx}-T_{yy})\\[5.0pt] {3(q_{x}+iq_{y})\over 2\sqrt{2}}&1+\sqrt{6}T_{xx}+\sqrt{6}T_{yy}&{3(q_{x}-iq_{y})\over 2\sqrt{2}}\\ \sqrt{3\over 2}(T_{xx}-T_{yy})&{3(q_{x}+iq_{y})\over 2\sqrt{2}}&1-{3q_{z}\over 2}+\sqrt{3\over 2}(T_{xx}+T_{yy}+2T_{zz})\\ \end{array}\right).$ (16)
Comparing the elements of the $\rho^{\psi}$ matrix, we can obtain $q_{x}=q_{y}=0,~{}q_{z}=-\mathcal{P}_{z},~{}T_{xx}={3P_{t}^{2}-1\over 2\sqrt{6}},~{}T_{yy}=-{1+3P_{t}^{2}\over 2\sqrt{6}},~{}T_{zz}=1/\sqrt{6}$ with other $T_{ij}(i\neq j)=0$. It can be seen that the tensor polarization of $\rho^{\gamma^{*}/\psi}$ comes from the spin correlation and transverse polarization of the initial lepton beams, with linear polarization being solely from the longitudinal polarization of beams. In the following paper, we focus solely on the transverse polarization of the initial lepton beams. The study of a longitudinally polarized electron beam has been recently addressed independently in other studies Salone et al. (2022); Zeng et al. (2023).
@@ II.0.2 $e^{+}e^{-}\to\gamma^{*}/\psi\to\Lambda(p\pi^{-})\bar{\Lambda}(\overline{p}\pi^{+})$
In the laboratory system, the decayed $\Lambda$ particle moves in the direction defined by polar and azimuthal angles $(\theta,\phi)$. We calculate the joint angular distribution in the $\Lambda$ helicity system, as illustrated in Fig. 1. In this system, the $z$-axis aligns with the $\Lambda$ particle’s direction of motion, the $y$-axis is perpendicular to the $\Lambda$ production plane, i.e., $\hat{y}=\hat{p}_{+}\times\hat{p}_{\Lambda}$, where $\hat{p}_{+}$ and $\hat{p}_{\Lambda}$ are the unit momentum vectors for the positron and $\Lambda$, respectively. The $x$-axis lies in the $\Lambda$ production plane, forming a right-handed $x$-$y$-$z$ coordinate system. After boosting the momenta of $p$ and $\pi^{-}$ in laboratory frame to the $\Lambda$ rest frame, they define the $\Lambda$ decay plane together with the momenta of $\Lambda$ in laboratory frame. The angle between the proton momentum in the $\Lambda$ decay plane and $\Lambda$ momentum in laboratory frame is defined as $\theta_{1}$, and the angle between the $\Lambda$ production and decay plane is defined as $\phi_{1}$. Similar helicity angles $(\theta_{2},\phi_{2})$ are defined in the same manner.
@@
$\rho_{1}^{i,j}(\theta,\phi)$$\equiv$ $\sum\limits_{k,k^{\prime}=\pm 1}\rho^{\gamma^{*}/\psi}_{k,k^{\prime}}{\cal D}_{k,i}^{1*}(\phi,\theta,0)\,{\cal D}^{1}_{k^{\prime},j}(\phi,\theta,0)$ (17)
@@
$+$ $\frac{3P_{T}^{2}}{3+\alpha_{\psi}}\cdot\left(\begin{array}[]{cccc}\alpha_{\psi}\sin^{2}\!\theta\cos\!2\phi&-{\beta_{\psi}{\sin\!\theta}}\sin\!2\phi&-{\beta_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi&0\\ -{\beta_{\psi}{\sin\!\theta}}\sin\!2\phi&(\alpha_{\psi}+\cos^{2}\!\theta)\cos\!2\phi&-(1+\alpha_{\psi})\cos\!\theta\sin\!2\phi&-{\gamma_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi\\ {\beta_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi&(1+\alpha_{\psi})\cos\!\theta\sin\!2\phi&(1+\alpha_{\psi}\cos^{2}\!\theta)\cos\!2\phi&-{\gamma_{\psi}{\sin\!\theta}}\sin\!2\phi\\ 0&{\gamma_{\psi}{\sin\!\theta\cos\!\theta}}\cos\!2\phi&-{\gamma_{\psi}{\sin\!\theta\sin\!2\phi}}&-\sin^{2}\!\theta\cos\!2\phi\\ \end{array}\right),$ (37)
@@
with vanishing longitudinal polarization component. Besides, the transverse polarization of beams introduces two new correlations $C_{xy}=-C_{yx}$, $C_{yz}=C_{zy}$ in addition to the four existing ones, which further constrain the decay parameters of hyperons, and thus provides more stringent tests of the CP violation.
@@
and Fig. 2 shows its three dimensional distribution. So that if the transverse polarization of the final-state particles is not measured, the effects of transverse polarizations are absent in the $\phi$-averaged cross section, albeit, present in the $\theta$-averaged cross section:
@@
with $(\theta_{0},\pi-\theta_{0})$ being the detector coverage of the solid angle around the interaction point (IP). Therefore the degree of transverse polarization would be measured through $e^{+}e^{-}\to\mu^{+}\mu^{-}$ and $e^{+}e^{-}$ azimuth angular distributions. The distribution of scattering angle has been previously explored in BES Bai et al. (1995), BESIII Ablikim et al. (2019b) and KEDR at the VEPP-4M Anashin et al. (2018), but the azimuth distributions did not receive any attention.
There are five global parameters to describe a process $e^{+}e^{-}\to B\overline{B}$ followed by single-step weak two-body decays of the hyperon $B$ and the antihyperon $\overline{B}$ Perotti et al. (2019). For decay $B\!\to\!b\pi$ and the corresponding charge conjugate (c.c.) decay mode $\overline{B}\!\to\!\overline{b}\overline{\pi}$, like $e^{+}e^{-}\to\gamma^{*}/\psi\to\Lambda(p\pi^{-})\overline{\Lambda}(\overline{p}\pi^{+})$, they are represented by the vector $\boldsymbol{\omega}\equiv(\alpha_{\psi},\Delta\Phi,\alpha_{-},{\alpha_{+}})$ with a priori known $P_{T}^{2}$ and $\alpha_{-}$ (or ${\alpha_{+}}$) being the decay parameter of $B\!\to\!b\pi$ (or $\overline{B}\!\to\!\overline{b}\overline{\pi}$). The joint angular distribution $\mathcal{W}(\boldsymbol{\xi})$ can be expressed with respect to the vector $\boldsymbol{\xi}\equiv(\Omega_{B},\Omega_{b},\Omega_{\overline{b}})$ representing a complete set of the kinematic variables describing a single-event configuration in the six dimensional phase space Hong et al. (2023); Zeng et al. (2023):
@@
Figure 3: The transverse polarization $P_{y}$ and $P_{x}$ of $\Lambda$ hyperon in all azimuthal angles in comparison of upper and lower spheres. The parameters of $e^{+}e^{-}\to\psi(3686)\to\Lambda\overline{\Lambda}$ with $\alpha_{\psi}=0.69$ and $\Delta\Phi=23^{\circ}$ Ablikim et al. (2023a) are used.
@@
the latter of which is corresponding to the number of events in the range of upper semi-sphere $45^{\circ}$ to $135^{\circ}$ and $225^{\circ}$ to $315^{\circ}$, and the former is the number of events in the remaining lower semi-sphere. Then $\mathcal{W}^{+}(\xi)-\mathcal{W}^{-}(\xi)$ is proportional to those terms of $P_{T}^{2}\cos 2\phi$ dependence and $\mathcal{W}^{+}(\xi)+\mathcal{W}^{-}(\xi)$ integrates out the $P_{T}^{2}$ terms which are of $\cos 2\phi$ modulation. To select the $\sin 2\phi$ modulation:
@@
the former of which is corresponding to the number of events in the range of upper semi-sphere $0^{\circ}$ to $90^{\circ}$ and $180^{\circ}$ to $270^{\circ}$, and the latter is the number of events in the remaining lower semi-sphere. Fig. 3 shows the transverse polarization $P_{y}$ and $P_{x}$ of baryon in all azimuthal angles in comparison of upper and lower spheres. Instead, one can investigate the moments of the joint angular distributions as shown in Sec. III.